# Principal Component Copulas for Capital Modeling and Systemic Risk

### v2

@@ 2.3 Copula density
Having discussed specific examples, we return to the general concept behind the copula construction technique, which is to generate novel dependence structures based on convolutions of analytically tractable probability distributions. Characteristic functions are convenient in obtaining analytic results for convolutions of stochastic variables. We summarize the most relevant properties of characteristic functions that we require in the following lemma. For clarity, we distinguish three different cases, where the first two situations are special versions of the more general third case. In the first case, all PCs $P_{j}$ are independent, so that all (univariate) characteristic functions $\phi_{P_{j}}$ multiply. In the second case and third case, we have that subgroups of generators ${\bf P}_{g}$ are independent, so that the corresponding (multivariate) characteristic functions $\phi_{{\bm{P}}_{g}}$ multiply. The PCs within such a subgroup are then not necessarily independent.
@@
Case 2. Let the characteristic function of the first copula generator $P_{1}$ and the higher copula generators ${\bf P}_{>1}$ be known and given by $\phi_{P_{1}}(t_{1})$ and $\phi_{\bm{P}_{>1}}(t_{2},...,t_{d})$. Then, we obtain for $\phi_{\bm{Y}}$ that:
$\phi_{\bm{Y}}(\bm{t})=E[e^{i\bm{t}^{\prime}\bm{Y}}]=E[e^{i\bm{t}^{\prime}\bm{WP}}]=E[e^{i(\bm{t}^{\prime}\bm{w_{1}})P_{1}}]E[^{i\bm{t}^{\prime}\bm{W}_{>1}\bm{P}_{>1}}]=\phi_{P_{1}}(\bm{w}^{\prime}_{1}\bm{t})\phi_{\bm{P}_{>1}}(\bm{W}^{\prime}_{>1}\bm{t}).$ (13)
Here, $\bm{w}_{1}$ is the first column of $\bm{W}$, while $\bm{W}_{>1}$ is the $d\times(d-1)$ matrix by omitting $\bm{w}_{1}$ from $\bm{W}$.
@@
$\phi_{\bm{Y}}(\bm{t})=E[e^{i\bm{t}^{\prime}\bm{Y}}]=E[e^{i\bm{t}^{\prime}\bm{WP}}]=\prod_{g=1}^{m}E[e^{i\bm{t}^{\prime}\bm{W}_{g}\bm{P}_{g}}]=\prod_{g=1}^{m}\phi_{{\bm{P}}_{g}}(\bm{W}^{\prime}_{g}\bm{t}).$ (14)
@@
This lemma therefore makes use of two convenient properties of characteristic functions, namely that the characteristic functions of (groups of) independent variables multiply, and that the marginal characteristic function is obtained from the multivariate characteristic function by setting all other arguments to zero. The first case of independent generators is relevant for the hyperbolic-normal PCC, because for the multivariate normal distribution it is equivalent to be uncorrelated and independent. The second case is relevant for the Skew $t_{1}$-$t_{d-1}$ PCC, because then all higher PCs are dependent. The third case allows general splitting of the PCs in subgroups and encompasses the first two cases. As a result, the multivariate distribution properties of the implicit copula for ${\bm{Y}}$ become accessible from the properties of the generator distributions. We can use the analytic results for the characteristic function to derive the marginal density and the marginal distribution for $Y_{i}$. These expressions are needed to determine the convolution copula density from Eq. (1) and for simulating from the copula. This is summarized in the following lemma that follows from the properties of the characteristic function and the Gil-Pelaez theorem:
@@
We start from Eq. (1) for the continuous probability density of an implicit copula. We consider the denominator in Eq. (1), for which we require the expressions $F_{Y_{i}}^{-1}$ and $f_{Y_{i}}$. Combining the results from Eqs. (14) and (15) with Eqs. (16) and (17), we obtain $F_{Y_{i}}$ and $f_{Y_{i}}$ as one-dimensional integrals over characteristic functions $\phi_{\bm{P}_{g}}$. Consequently, $F_{Y_{i}}^{-1}$ is obtained by one-dimensional inversion. Next, we proceed with the expression for the numerator. We have that $\bm{P}=\bm{W}^{\prime}\bm{Y}$ and that $\bm{P}_{g}=\bm{W}_{g}^{\prime}\bm{Y}$ with $\bm{W}_{g}$ the $d\times d_{g}$ matrix containing the columns $j$ from $\bm{W}$ belonging to $P_{j}$ in subgroup $g$. As a result, we have:
@@
For a factor copula based on multiple factors with general distributions, we have that Eq. (18) does not apply, because it is based on a latent factor structure, rather than an orthogonal transformation. As a result, determining $f_{\mathbf{Y}}$ and $c(u_{1},...,u_{d})$ requires integrating over common factors in that case.
We can specify the PCC ($C_{\bm{\Lambda},\mathbf{W},\bm{\alpha}}$) by the eigenvalues $\Lambda_{j}$ which describe the variance of the PCs, by the orthogonal matrix elements $w_{i,j}$ describing the linear transformation from $P_{j}$ to $Y_{i}$, and by possible shape parameters $\bm{\alpha}$, which allow $P_{j}$ to be skewed or heavy-tailed. By definition of the PCC, we require the parameters $\Lambda_{j}$ and $w_{i,j}$ to form a valid correlation matrix, since they satisfy $\bm{\rho}_{\bm{Y}}=\bm{W}\bm{\Lambda}\bm{W}^{\prime}$. This leads to restrictions on admissible parameter values for $\Lambda_{j}$ and $w_{i,j}$, which can be cumbersome to keep track of while performing estimation in high dimensions. As a result, we prefer to parametrize the PCC by the correlation matrix $\bm{\rho}_{\mathbf{Y}}$ and the shape parameters $\bm{\alpha}$. In high-dimensional applications, we prefer to estimate $\bm{\rho}_{\mathbf{Y}}$ with a moment estimator, such that it is automatically a valid correlation matrix, after which we perform the decomposition in eigenvalues and eigenvectors. If the eigenvalues $\Lambda_{j}$ are non-degenerate, the eigenvalue decomposition of $\bm{\rho}_{\mathbf{Y}}$ is unique up to the sign for the eigenvectors. In practice, it is common to enforce a sign convention, for example, that the largest eigenvector weight is taken to be positive.
We note that the PCC is not equivalent to merely specifying a copula between PCs, since this is not sufficient to fully specify the copula between risk factors $Y_{i}$. For example, an independence copula between PCs can lead to strong tail dependence between risk factors, when the PCs have fatter tails than Gaussian (see Proposition 1). The PCC fully specifies the implied copula density for the risk factors $Y_{i}$ that is generated by the PCs $P_{j}$. To the best of our knowledge, the implicit copula that is generated by PCA models has not been systematically studied before.
@@ 3 Estimation procedures
Having introduced the PCC and considered several properties, we next discuss the estimation procedures. We assume that a $d$-dimensional data sample consisting of $n$ iid observations $\bm{X}_{t}$ with $t=1,...,n$ has been observed.
@@ 3.1 Data exploration and model initiation
We start with explaining a procedure to explore the data sample, which is also convenient to initialize estimation procedures. We follow the scheme of Fig. 1a, which amounts to performing PCA on normal scores. The advantage of this scheme is that the first three steps can be performed directly on the original data $\bm{X}_{t}$ without requiring knowledge about the parameters of the implicit copula. Since we are only interested in the copula for $\bm{X}_{t}$, we transform the original data $X_{i,t}$ in the first step of Fig. 1a to pseudo copula observations using $U_{i,t}={\rm rank}(X_{i,t})/(n+1)$. The second step of Figure 1a is also readily performed by obtaining the normal scores $Y_{i,t}=\Phi^{-1}(U_{i,t})$. As a result, the data is symmetrized and outliers are reduced before applying PCA. The third step is to apply PCA by determining the eigenvalues and eigenvectors of the sample correlation matrix of $\bm{Y}$ and decompose this matrix in the diagonal eigenvalue matrix $\bm{\Lambda}$ and the orthogonal matrix $\bm{W}$. Using $\bm{P}_{t}=\bm{W}^{\prime}\bm{Y}_{t}$ we obtain an initial approximation of the PC observations $P_{j,t}$ based on matrix multiplication of normal scores, which is easily performed.
@@ 3.2 Maximum likelihood estimation
$\ell(\bm{\theta};U_{1,t},..,U_{d,t})$ $=$ $\sum_{t=1}^{n}{\rm ln}\,c(U_{1,t},...,U_{d,t};\bm{\theta})=\sum_{t=1}^{n}{\rm ln}\,f_{\mathbf{Y}}(Y_{1,t},...,Y_{d,t};\bm{\theta})-\sum_{t=1,i=1}^{n,d}{\rm ln}\,f_{Y_{i}}(Y_{i};\bm{\theta}).$ (19)
Here, $Y_{i,t}=F^{-1}_{Y_{i}}(U_{i,t};\bm{\theta})$, while $f_{Y_{i}}$, $F_{Y_{i}}$ and $f_{\mathbf{Y}}$ are given by Eqs. (16), (17) and (18). The first two equations require the marginal characteristic function that is obtained from Eqs. (14) and (15). The maximum likelihood estimator is given by:
@@
MLEs based on pseudo copula observations have the desirable asymptotic properties of being consistent, asymptotically normal and efficient, as shown by Genest et al. [26]. The main challenge for the PCC is then to find a fast and accurate way to determine the copula density and the loglikelihood in Eq. (19), which can consequently be used in optimization algorithms for Eq. (20).
@@ 3.2.1 Numerical approach
Because $Y_{i}$ has unit variance and because the COS method converges exponentially, we typically find accurate results with the following choices: $a=-10,b=10$ and $N_{c}=100$. Having implemented the COS method, we can efficiently evaluate $f_{Y_{i}}(y_{i})$ and $F_{Y_{i}}(y_{i})$ on a coordinate grid of our choice. If we evaluate $F_{Y_{i}}(y_{i})=u_{i}$ on a sufficiently fine grid, we immediately also obtain $F^{-1}_{Y_{i}}(u_{i})=y_{i}$ on the same grid. Any other value for $F^{-1}_{Y_{i}}(u_{i})$ can be obtained by interpolation. An alternative method to sample from the inverse distribution function is stochastic collocation [27].
Next, the optimization of Eq. (20) is performed to obtain the ML estimates. The optimization is initialized with the linear correlation matrix of the normal scores $\Phi^{-1}(u_{i,t})$, which is the ML estimator for the correlation parameters of the Gaussian copula, and with plausible starting values of the shape parameters $\bm{\alpha}$. Gradient descent is then employed to update the parameters $\bm{\rho}_{\mathbf{Y}}$ and $\bm{\alpha}$. After each update of $\bm{\rho}_{\mathbf{Y}}$, we reperform eigenvalue decomposition (PCA) to obtain $\bm{\Lambda}$ and $\mathbf{W}$. The correlation matrices are kept positive semi-definite by requiring the eigenvalues of $\bm{\Lambda}$ to be nonnegative. An alternative method would be to parameterise the correlation matrix in a convenient way [28].
@@ 3.3 Generalized Method of Moments
Besides ML, another popular class of estimators is the generalized method of moments (GMM) [29]. It starts from a vector-valued function $\bm{g}(\bm{Y},\bm{\theta})$ with an expectation value of zero at the true parameter values $\bm{\theta}_{0}$. The expectation value of $\bm{g}$ can be estimated with the vector valued sample mean:
@@ 3.3.1 Numerical approach
Next, we explain how we numerically arrive at the solution of the GMM equations following from the conditions in Eq. (26). Note that $F_{Y_{i}}^{-1}$ depends itself on $\bm{\rho}_{\mathbf{Y}}$ and $\bm{\alpha}$, making the corresponding equations nonlinear. However, in the applications of our interest, the dependence is weak, so that an iterative procedure can be used. The iterative procedure initializes the correlation matrix $\hat{\rho}^{(0)}_{\bm{Y}}$ with the sample correlation of the normal scores $\Phi^{-1}(u_{i,t})$. Next, we decompose the correlation matrix into eigenvectors using PCA. In general, the following holds for the PCC:
@@
An advantage of the iterative approach is that it allows the correlation matrix to be determined with a moment estimator. This avoids having to numerically optimize over correlation parameters, which is required for ML or Simulated Method of Moments when there is no closed-form expression for the correlation parameters in terms of the data available [8]. As a result, these methods have been mainly applied to high-dimensional copulas with a reduced number of parameters by imposing additional structure [8, 9]. Although parsimonious model structures are favorable, it is also beneficial to have a technique that avoids the necessity to make a priori assumptions on the correlation structure. In that case, a reduction in parameters can still be imposed, but only when it is deemed necessary in terms of estimation stability and accuracy.
In the next section, we show that the algorithm performs very well in detecting PCs with large eigenvalues that can have a nontrivial structure based on their shape parameters. In the studied examples, the first eigenvalues and eigenvectors are found to be robust against changes in the shape parameters. Consequently, the initialization steps already give accurate results, and only a few iteration steps are required to obtain the full solution. In the next section, we show very good performance with the hybrid estimator of Algorithm 1 for large copula models. The performance remains intact in high dimensions, which is considered an attractive feature of the algorithm.

### v3

@@ 2.3 Copula density
Having discussed specific examples, we return to the general concept behind the copula construction technique, which is to generate novel dependence structures based on convolutions of analytically tractable probability distributions. In Fig. 2, we illustrate the Data Generating Process of the Principal Component Copula. Characteristic functions are convenient in obtaining analytic results for convolutions of stochastic variables. We summarize the most relevant properties of characteristic functions that we require in the following lemma. For clarity, we distinguish three different cases, where the first two situations are special versions of the more general third case. In the first case, all PCs $P_{j}$ are independent, so that all (univariate) characteristic functions $\phi_{P_{j}}$ multiply. In the second case and third case, we have that subgroups of generators ${\bf P}_{g}$ are independent, so that the corresponding (multivariate) characteristic functions $\phi_{{\bm{P}}_{g}}$ multiply. The PCs within such a subgroup are then not necessarily independent.
Figure 2: Visualization of the data generating process for the Principal Component Copula (PCC). The joint distribution of uncorrelated generators $P_{j}$ forms the basis of the implicit copula. The orthogonal matrix $\bm{W}$ performs a rotation to introduce correlations between stochastic variables $Y_{i}$. Each marginal distribution $F_{Y_{i}}$ can be computed using convolution to complete the copula construction for $U_{i}$. Tail dependence is generated in a flexible and controlled way by distribution choices in the parallel and orthogonal directions.
@@
Case 2. Let the characteristic function of the first copula generator $P_{1}$ and the higher copula generators ${\bf P}_{>}$ be known and given by $\phi_{P_{1}}(t_{1})$ and $\phi_{\bm{P}_{>}}(t_{2},...,t_{d})$. Then, we obtain for $\phi_{\bm{Y}}$ that:
$\phi_{\bm{Y}}(\bm{t})=E[e^{i\bm{t}^{\prime}\bm{WP}}]=E[e^{i(\bm{t}^{\prime}\bm{w_{1}})P_{1}}]E[^{i\bm{t}^{\prime}\bm{W}_{>}\bm{P}_{>}}]=\phi_{P_{1}}(\bm{w}^{\prime}_{1}\bm{t})\phi_{\bm{P}_{>}}(\bm{W}^{\prime}_{>}\bm{t}).$
Here, $\bm{w}_{1}$ is the first column of $\bm{W}$, while $\bm{W}_{>}$ is the $d\times(d-1)$ matrix by omitting $\bm{w}_{1}$ from $\bm{W}$.
@@
$\phi_{\bm{Y}}(\bm{t})=E[e^{i\bm{t}^{\prime}\bm{WP}}]=\prod_{g=1}^{m}E[e^{i\bm{t}^{\prime}\bm{W}_{g}\bm{P}_{g}}]=\prod_{g=1}^{m}\phi_{{\bm{P}}_{g}}(\bm{W}^{\prime}_{g}\bm{t}).$ (13)
@@
This lemma therefore makes use of two convenient properties of characteristic functions, namely that the characteristic functions of (groups of) independent variables multiply, and that the marginal characteristic function is obtained from the multivariate characteristic function by setting all other arguments to zero. The first case of independent generators is relevant for the hyperbolic-normal PCC, because for the multivariate normal distribution it is equivalent to be uncorrelated and independent. The second case is relevant for the Skew $t_{1}$-$t_{d-1}$ PCC, because then all higher PCs are dependent. The third case allows general splitting of the PCs in subgroups and encompasses the first two cases. As a result, the multivariate distribution properties of the implicit copula for ${\bm{Y}}$ become accessible from the properties of the generator distributions. We can use the analytic results for the characteristic function to derive the marginal density and the marginal distribution function for $Y_{i}$. These expressions are needed to determine the convolution copula density from Eq. (1) and for simulating from the copula. This is summarized in the following lemma that follows from the properties of the characteristic function and the Gil-Pelaez theorem:
@@
We start from Eq. (1) for the continuous probability density of an implicit copula. We consider the denominator in Eq. (1), for which we require the expressions $F_{Y_{i}}^{-1}$ and $f_{Y_{i}}$. Combining the results from Eqs. (13) and (14) with Eqs. (15) and (16), we obtain $F_{Y_{i}}$ and $f_{Y_{i}}$ as one-dimensional integrals over characteristic functions $\phi_{\bm{P}_{g}}$. Consequently, $F_{Y_{i}}^{-1}$ can be obtained by one-dimensional inversion. Next, we proceed with the expression for the numerator. We have that $\bm{P}=\bm{W}^{\prime}\bm{Y}$ and that $\bm{P}_{g}=\bm{W}_{g}^{\prime}\bm{Y}$ with $\bm{W}_{g}$ the $d\times d_{g}$ matrix containing the columns $j$ from $\bm{W}$ belonging to $P_{j}$ in subgroup $g$. As a result, we have:
@@
For a factor copula based on multiple factors, we have that Eq. (17) does not apply, because it is based on a latent factor structure [8], rather than an orthogonal transformation. As a result, determining $f_{\mathbf{Y}}$ and $c(u_{1},...,u_{d})$ requires integrating over all common factors in that case.
We can specify the PCC ($C_{\bm{\Lambda},\mathbf{W},\bm{\alpha}}$) by the eigenvalues $\Lambda_{j}$ which describe the variance of the PCs, by the orthogonal matrix elements $w_{i,j}$ describing the linear transformation from $P_{j}$ to $Y_{i}$, and by the shape parameters $\bm{\alpha}$, which allow $P_{j}$ to be skewed or heavy-tailed. By definition of the PCC, we require the parameters $\Lambda_{j}$ and $w_{i,j}$ to form a valid correlation matrix, since they satisfy $\bm{\rho}_{\bm{Y}}=\bm{W}\bm{\Lambda}\bm{W}^{\prime}$. This leads to restrictions on admissible parameter values for $\Lambda_{j}$ and $w_{i,j}$, which can be cumbersome to keep track of while performing estimation in high dimensions. As a result, we prefer to parametrize the PCC by the correlation matrix $\bm{\rho}_{\mathbf{Y}}$ and the shape parameters $\bm{\alpha}$. In high-dimensional applications, we prefer to estimate $\bm{\rho}_{\mathbf{Y}}$ with a moment estimator, such that it is automatically a valid correlation matrix, after which we perform the decomposition in eigenvalues and eigenvectors. If the eigenvalues $\Lambda_{j}$ are non-degenerate, the eigenvalue decomposition of $\bm{\rho}_{\mathbf{Y}}$ is unique up to the sign for the eigenvectors. In practice, it is common to enforce a sign convention, for example, that the largest eigenvector weight is taken to be positive.
We note that the PCC is not equivalent to merely specifying a copula between PCs, since this is not sufficient to fully specify the copula for the stochastic variables $Y_{i}$. For example, an independence copula between PCs can lead to strong tail dependence between risk factors, when the PCs have fatter tails than Gaussian (see Proposition 1). The PCC fully specifies the implied copula density for the stochastic variables $Y_{i}$ that is generated by the PCs $P_{j}$. To the best of our knowledge, the implicit copula that is generated by PCA has not been systematically studied before.
@@ 3 Estimation procedures
Having introduced the PCC and considered several properties, we next discuss the estimation procedures. In case of financial returns, it is common to first apply a time series filter (e.g. GARCH) to arrive at devolatized returns, after which the dependence structure is studied. In Section 4.2.1, we give a concrete example for such a time series model. In this section, we assume that a $d$-dimensional data sample is observed that consists of $n$ iid observations $\bm{X}_{t}$ with $t=1,...,n$.
@@ 3.1 Data exploration and model initiation
We start with explaining a procedure to explore the data sample, which is also convenient to initialize estimation procedures. We follow the scheme of Fig. 3a, which amounts to performing PCA on normal scores. The advantage of this scheme is that the first three steps can be performed directly on the original data $\bm{X}_{t}$ without requiring knowledge about the parameters of the implicit copula. Since we are only interested in the copula for $\bm{X}_{t}$, we transform the original data $X_{i,t}$ in the first step of Fig. 3a to (pseudo) copula observations using $U_{i,t}={\rm rank}(X_{i,t})/(n+1)$. The second step of Figure 3a is also readily performed by obtaining the normal scores $Y_{i,t}=\Phi^{-1}(U_{i,t})$. As a result, the data is symmetrized and outliers are reduced before applying PCA. The third step is to apply PCA by determining the eigenvalues and eigenvectors of the sample correlation matrix of $\bm{Y}$ and decompose this matrix in the diagonal eigenvalue matrix $\bm{\Lambda}$ and the orthogonal matrix $\bm{W}$. Using $\bm{P}_{t}=\bm{W}^{\prime}\bm{Y}_{t}$ we obtain an initial approximation of the PC observations $P_{j,t}$ based on matrix multiplication of normal scores, which is easily performed.
@@ 3.2 Maximum likelihood estimation
$\ell(\bm{\theta};U_{1,t},..,U_{d,t})$ $=$ $\sum_{t=1}^{n}{\rm ln}\,c(U_{1,t},...,U_{d,t};\bm{\theta})$ (18)
$=$ $\sum_{t=1}^{n}{\rm ln}\,f_{\mathbf{Y}}(Y_{1,t},...,Y_{d,t};\bm{\theta})-\sum_{t=1,i=1}^{n,d}{\rm ln}\,f_{Y_{i}}(Y_{i};\bm{\theta}).$
Here, $Y_{i,t}=F^{-1}_{Y_{i}}(U_{i,t};\bm{\theta})$, while $f_{Y_{i}}$, $F_{Y_{i}}$ and $f_{\mathbf{Y}}$ are given by Eqs. (15), (16) and (17). The first two equations require the marginal characteristic function that is obtained from Eqs. (13) and (14). The MLE is given by:
@@
MLEs based on pseudo copula observations have the desirable asymptotic properties of being consistent, asymptotically normal and efficient, as shown by [26]. The main challenge for the PCC is then to find a fast and accurate way to determine the copula density and the loglikelihood in Eq. (18), which can consequently be used in optimization algorithms for Eq. (19).
@@ 3.2.1 Numerical approach
Because $Y_{i}$ has unit variance and because the COS method converges exponentially, we typically find accurate results using: $a=-10,b=10$ and $N_{c}=100$. Having implemented the COS method, we can efficiently evaluate $f_{Y_{i}}(y_{i})$ and $F_{Y_{i}}(y_{i})$ on a coordinate grid of our choice. If we evaluate $F_{Y_{i}}(y_{i})=u_{i}$ on a sufficiently fine grid, we immediately also obtain $F^{-1}_{Y_{i}}(u_{i})=y_{i}$ on the same grid. Any other value for $F^{-1}_{Y_{i}}(u_{i})$ can be obtained by interpolation. An alternative method to sample from the inverse distribution function is stochastic collocation [27].
Next, the optimization of Eq. (19) is performed to obtain the ML estimates. The optimization is initialized with the linear correlation matrix of the normal scores $\Phi^{-1}(U_{i,t})$, which is the ML estimator for the correlation parameters of the Gaussian copula, and with plausible starting values of the shape parameters $\bm{\alpha}$. Gradient descent is then employed to update the parameters $\bm{\rho}_{\mathbf{Y}}$ and $\bm{\alpha}$. After each update of $\bm{\rho}_{\mathbf{Y}}$, we reperform eigenvalue decomposition (PCA) to obtain $\bm{\Lambda}$ and $\mathbf{W}$. The correlation matrices are kept positive semi-definite by requiring the eigenvalues of $\bm{\Lambda}$ to be nonnegative.
@@ 3.3 Generalized Method of Moments
Besides ML, another popular class of estimators is the generalized method of moments (GMM) by [28]. It starts from a vector-valued function $\bm{g}(\bm{Y},\bm{\theta})$ with an expectation value of zero at the true parameter values $\bm{\theta}_{0}$. The expectation value of $\bm{g}$ can be estimated with the vector valued sample mean:
@@ 3.3.1 Numerical approach
Next, we explain how we numerically arrive at the solution of the GMM equations following from the conditions in Eq. (25). Note that $F_{Y_{i}}^{-1}$ depends itself on $\bm{\rho}_{\mathbf{Y}}$ and $\bm{\alpha}$, making the corresponding equations nonlinear. We propose an iterative procedure to solve the equations. The procedure initializes the correlation matrix $\hat{\rho}^{(0)}_{\bm{Y}}$ with the sample correlation of the normal scores $\Phi^{-1}(u_{i,t})$. Next, we decompose the correlation matrix into eigenvectors using PCA. In general, the following holds for the PCC:
@@
An advantage of the iterative approach is that it allows the correlation matrix to be determined with a moment estimator. This avoids having to numerically optimize over many correlation parameters, which is required for ML or Simulated Method of Moments when there is no closed-form expression for the correlation parameters in terms of the data available [8]. As a result, these methods have been mainly applied to high-dimensional copulas with a reduced number of parameters by imposing additional structure to the correlation matrix. Although reducing complexity can be favorable, it is also beneficial to have a technique that avoids the necessity to make a priori assumptions on the correlation structure. In that case, a reduction in parameters can still be imposed, but only when it is deemed necessary in terms of estimation stability and accuracy.
In the next section, we show that the algorithm performs well in detecting PCs with large eigenvalues. In the studied examples, the estimation of the first eigenvalues and eigenvectors are found to be robust to the changes in shape parameters of the PC distributions. Consequently, the initialization steps already give accurate results, and only a few iteration steps are required to obtain the full solution. In the next section, we show good performance with the hybrid estimator of Algorithm 1 for large copula models in high dimensions.